OpenStudy (anonymous):

how to find the domain ? *g(x)=square root of x over 2x^2 + x -1

OpenStudy (anonymous):

help please :))

OpenStudy (anonymous):

sqrt requires positive x. The denominator can take all positive x except x=1/2 (division by 0).

OpenStudy (anonymous):

It's the square root of x or the square root of everything?

OpenStudy (anonymous):

\[\sqrt{x} \over 2x ^{2}+x-1\]

OpenStudy (anonymous):

so you have to make sure that \[x\geq 0\] and also that \[2x^2+x-1\neq 0\]

OpenStudy (anonymous):

oh right, what estudier said